OpenStudy (anonymous):

How do you solve 3x² - 12x + 12 = 0 ??

OpenStudy (anonymous):

Xsquared - 2x -2x +4 x (x-2) -2(x-2)=0 (x-2)(x-2)=0 x=2

OpenStudy (anonymous):

Yes, when you factor it you should get: 3(x^2 - 4x + 4)=0. Using FOIL method, we can re-write the equation to 3(x-2)(x-2)=0 thus the solution would be x=2 as a double root

OpenStudy (anonymous):

factor out the 3 3*(x^2-4x+4)=0 3*(x-2)^2

OpenStudy (dumbcow):

deja vu ... i know i've just seen this exact problem a little while ago

OpenStudy (anonymous):

lol @dumbcow me i asked the same question

OpenStudy (dumbcow):

haha nice

OpenStudy (dumbcow):

and now you are answering it...thats awesome, that means you get it now

OpenStudy (anonymous):

You should not need to use the quadratic formula. 3x² - 12x + 12 = 0 [multiply 3 from x^2 by +12] = 36 [Find factors of 36 that sum to 12] -6 x -6 = 12 and [-6] + [-6] = -12 Substitute into -12x 3x² - 6x - 6x + 12 = 0 [factorise in parts] (3x² - 6x) - (6x + 12) = 0 3x(x - 2) - 6(x - 2) = 0 Answer 3(x - 2)(x - 2) Hence x = 2